hOw to calculate the number Of nuetrons present in 9mg of O18 ..please give proper answer

Grade:12

1 Answers

Mohd Mujtaba
6 years ago
Here no of neutron in one atom of 18O=18-8=10.Therefore no of neutron in 18 g of 18O=10×Na.So no of neutron in 9mg of O18=(10×Na×9×10^-3)÷18=3×10^21 So answer is 3×10^21